Derive schrodinger wave equation in its time dependent form


1
Expert's answer
2022-02-10T15:05:27-0500

"i\hbar\frac{\partial}{\partial t}\psi=\hat H(p,q)\psi,"

"\hat H=-\frac{\hbar^2}{2m}\nabla^2+V(\vec r,t),"

"i\hbar\frac{\partial}{\partial t}\psi(\vec r,t)=[-\frac{\hbar^2}{2m}\nabla^2+V(\vec r,t)]\psi(\vec r,t)."